# Sums related to Euler’s totient function

### v5

@@ 1. Introduction
Let $b_{1},\ldots,b_{k}$ be distinct integers, $a\in\mathbb{N}$, $s\in\mathbb{N}$. Let $k\geq 2,$ $x\geq 3$, $(\ln x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\log x$ for all $1\leq i\leq k$. We set
@@ 2. Notation
$S=\sum_{b\in\Omega}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}=\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}S_{0}.$ (4.3)
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\log x$ for all $1\leq i\leq k$, and $|f(b)|\leq(2\log x)^{k}$. Suppose that $k\geq\log\log x$. Since $n/\varphi(n)\leq c\log\log(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
@@
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We can assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
@@
$\omega(n)=\#\{p\leq x:f(p)\equiv 0\text{ (mod $n$)}=\sum_{i=1}^{l}\pi(x;n,\alpha_{i}).$ (5.2)

### v6

@@ 1. Introduction
Let $b_{1},\ldots,b_{k}$ be distinct integers, $a\in\mathbb{N}$, $s\in\mathbb{N}$. Let $k\geq 2,$ $x\geq 3$, $(\log x)^{-9/10}\leq\eta\leq 1$, and $|b_{i}|\leq\log x$ for all $1\leq i\leq k$. We set
@@ 2. Notation
$S=\sum_{b\in\Omega}\Big{(}\frac{a^{k+1}|f(b)|}{\varphi(a^{k+1}|f(b)|)}\Big{)}^{s}\leq\Big{(}\frac{a}{\varphi(a)}\Big{)}^{s}\sum_{b\in\Omega}\Big{(}\frac{|f(b)|}{\varphi(|f(b)|)}\Big{)}^{s}.$ (4.3)
Let $b\in\Omega$. Hence, $|b-b_{i}|\leq 2\log x$ for all $1\leq i\leq k$, and therefore $|f(b)|\leq(2\log x)^{k}$. Suppose that $k\geq\log\log x$. Since $n/\varphi(n)\leq c\log\log(n+2)$ for any positive integer $n$, where $c$ is an absolute positive constant, we obtain
@@
We set $a_{p}=f(p)$. It is clear that $N=\pi(x)$. There exists a positive constant $\tau$ depending only on $f$ such that $f(n)\leq\tau n^{d}$ for any positive integer $n$. We take $M=\tau x^{d}$ and $\alpha=1/2$. We assume that $x\geq x_{0}$, where $x_{0}>0$ is a large constant depending only on $f$. By (4.1) and (4.2) we have $y=(\log M)^{\alpha}\leq(\log x)^{3/4}$ and $n\leq\sqrt{x}$ for any $n\in\mathcal{D}$.
@@
$\omega(n)=\#\{p\leq x:f(p)\equiv 0\text{ (mod $n$)}\}=\sum_{i=1}^{l}\pi(x;n,\alpha_{i}).$ (5.2)